OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (triciaal):

let x = 2 and put in g(x) when you get that value use that as x in f(x)

OpenStudy (unklerhaukus):

\[g(x)=x^2-6x-7\]\[f(x)=x+8\] \[f(g(x))=f(x^2-6x-7)=(x^2-6x-7)+8\] \[f(g(2))=f((2)^2-6(2)-7)=\big((2)^2-6(2)-7\big)+8\]